OpenStudy (anonymous):

x(x^2-1)+2-[x-(3-x^2)x]

OpenStudy (anonymous):

distributive property \[x(x^2-1)+2-(x-(3-x^2)x)\] \[=x^3-x+2-(x-(3x-x^3))\] \[=x^3-x+2-(x-3x+x^3)\] \[=x^3-x+2-x+3x-x^3\] and now there is very little left after you combine like terms

OpenStudy (anonymous):

you are left only with \(x+2\)

OpenStudy (anonymous):

can u explain further explain how did u come up with that answer?